# Direct in-situ observations of wave-induced floe collisions in the deeper Marginal Ice Zone

### v1

@@ 5 Conclusion
We recovered raw data time series of wave-in-ice motion from a buoy that got stranded on the northern Icelandic coast. Statistically significant extreme residual acceleration events, i.e., spikes in the acceleration signal once the incoming wave motion is subtracted, are detected with a statistical confidence of 99% in four of the raw data files. Of these four, two are then found to have clear signs of collisions in the raw data. These manifest as extreme residual acceleration events corresponding to spikes in the horizontal acceleration measured by the buoy. The direction of these spikes is generally aligned with the dominating wave direction incoming from the open ocean that is predicted by ocean models, and corresponds well to the local wave direction observed in the horizontal in-situ acceleration time series. Moreover, the occurrence of the acceleration events is phase locked to the incoming wave signal.
Based on these evidence, we conclude that the data we recovered contain wave-induced floe-floe collisions between the floe on which the buoy was sitting, and neighboring floes. This is the first time, to our knowledge, that such clear floe-floe collision evidences are provided deep into the Marginal Ice Zone (MIZ). This proves that collisions between adjacent ice floes can happen deep in the MIZ, and that confirms that floe-floe collisions are a relevant mechanism that can contribute significantly to the wave energy dissipation and the associated increased water turbulence levels, including in the deep MIZ.

### v3

@@ 5 Conclusion
We recovered raw data time series of wave-in-ice motion from a buoy that got stranded on the northern Icelandic coast. Statistically significant extreme residual acceleration events, i.e., spikes in the acceleration signal once the incoming wave motion is subtracted, are detected with a statistical confidence of 99% in four of the 20 minute long raw data files. Of these four, two are then found to have clear signs of collisions in the raw data. These manifest as extreme residual acceleration events corresponding to spikes in the horizontal acceleration measured by the buoy. The direction of these spikes is generally aligned with the dominating wave direction incoming from the open ocean that is predicted by ocean models, and corresponds well to the local wave direction observed in the horizontal in-situ acceleration time series. Moreover, the occurrence of the acceleration events is phase locked to the incoming wave signal.
Based on this evidence, we conclude that the data we recovered contain wave-induced floe-floe collisions between the floe on which the buoy was sitting, and neighboring floes. This is the first time, to our knowledge, that such clear floe-floe collision evidence are provided deep into the Marginal Ice Zone (MIZ). This proves that collisions between adjacent ice floes can happen deep in the MIZ, and that confirms that floe-floe collisions are a relevant mechanism that can contribute significantly to the wave energy dissipation and the associated increased water turbulence levels, including in the deep MIZ. Based on these evidence, we conclude that the data we recovered contain wave-induced floe-floe collisions between the floe on which the buoy was sitting, and neighboring floes. This is the first time, to our knowledge, that such clear floe-floe collision evidences are provided deep into the Marginal Ice Zone (MIZ). This proves that collisions between adjacent ice floes can happen deep in the MIZ, and that confirms that floe-floe collisions are a relevant mechanism that can contribute significantly to the wave energy dissipation and the associated increased water turbulence levels, including in the deep MIZ.